### v1

# Few-Shot Multimodal Medical Imaging: A Theoretical Framework

Abstract
Medical imaging relies heavily on large, labeled datasets. But, unfortunately, they are not always easily accessible in clinical settings. Additionally, many practitioners often face various structural obstacles like limited data availability, fragmented data systems, and unbalanced datasets. These barriers often lead to the increased diagnostic uncertainty, underrepresentation of certain conditions, reduced model robustness, and biased diagnostic decisions. In response to these challenges, approaches such as transfer learning, meta-learning, and multimodal fusion have made great strides. However, they still need a solid theoretical justification for why they succeed or fail in situations where data is scarce. To address this gap, we propose a unified theoretical framework that characterizes learning and inference under low-resource medical imaging conditions. We first formalize the learning objective under few-shot conditions and compute sample complexity constraints to estimate the smallest quantity of data needed to achieve clinically reliable accuracy. Then based on ideas from PAC-learning and PAC-Bayesian theory, we explain how multimodal integration encourages generalization and quantifies uncertainty under sparse supervision. We further propose a formal metric for explanation stability, offering interpretability guarantees under low-data conditions. Taken together, the proposed framework establishes a principled foundation for constructing dependable, data-efficient diagnostic systems by jointly characterizing sample efficiency, uncertainty quantification, and interpretability in a unified theoretical setting.
I Introduction
As an essential pillar of modern healthcare, medical imaging underpins diagnosis, therapeutic decision-making, and longitudinal disease monitoring [1]. It is especially evident in identifying rare diseases in low-resource healthcare systems, which lack substantial, well-annotated datasets. Also, single-modality images being used in these systems frequently yield diagnostic information which are not enough. Most of the earlier diagnostic modeling approaches relied on single data modalities, which limited their ability to grasp the full extent of complementary clinical cues [2]. For mitigating this limitation, multimodal imaging has emerged as a promising approach, where information from multiple sources is combined to produce richer, condensed as well as more informative representations [3, 4]. Although these integrated methods have improved diagnostic precision, their success is also often constrained by the scarcity of labeled data, especially for rare diseases.
Few-shot learning (FSL) has emerged as a promising approach to address this challenge. FSL allows models to generalize from few labeled samples [5], and meta-learning improves FSL by enabling quick adaptation to new tasks with less supervision [6]. However, there remains no clear theoretical understanding of how much data is sufficient, how uncertainty behaves under restricted supervision, or how interpretability can be preserved. Motivated by this gap, we develop a theoretical framework for low-resource medical imaging, grounded in Vapnik–Chervonenkis (VC) and Probably Approximately Correct (PAC) learning theories, to formalize the relationships among sample complexity, uncertainty, and interpretability, and to introduce a new metric—explanation variance—for assessing interpretability stability under data scarcity.
II Background and Related Work
II-A Low-Resource and Few-Shot Medical Imaging
One of the central challenges in medical imaging is building reliable models when labeled data are limited. Few-shot learning (FSL) provides a way to train models that can generalize from a small number of annotated examples [7]. It has become an essential approach for overcoming the persistent shortage of annotated medical images and the limited size of publicly available datasets [8, 9]. Although recent studies have explored various few-shot and task-adaptive methods, most lack theoretical grounding in how much labeled data are actually required to achieve clinically dependable performance. This gap has led to increasing efforts to develop formal frameworks that characterize the sample complexity of medical imaging models.
II-B Multimodal Integration and Information Gain
Multimodal techniques are increasingly adopted in both healthcare research and clinical practice because they combine diverse sources of data to produce systems that are more adaptive, reliable, and context-aware [10, 11]. This shift reflects a move from traditional, centralized approaches toward personalized, patient-centered models of care. However, much of the progress in multimodal learning remains empirical, with limited theoretical insight into why and how multimodal systems outperform single-modality counterparts. Furthermore, few studies have examined how interactions among modalities influence learnability and generalization. Establishing a rigorous information-theoretic foundation for multimodal learning is therefore essential to guide model design and interpretation in medical applications.
II-C Uncertainty Quantification under Sparse Supervision
In clinical research and practice, accurately estimating uncertainty is essential, as predictive errors can have serious consequences. Uncertainty quantification (UQ) methods assess the reliability of predictive models and support safer decision-making [11, 12]. Techniques such as variational inference [15, 16], Monte Carlo dropout [14], approximate Bayesian inference [17], and Bayesian deep ensembles [18] have been widely employed to enhance model robustness and reliability across diverse domains [13]. In healthcare, uncertainty estimation is particularly valuable for detecting anomalous cases, flagging atypical or ambiguous results, and strengthening clinician confidence in computational assessments [19].
Despite these advancements, the majority of uncertainty quantification (UQ) methodologies continue to be predominantly empirical, and theoretical comprehension of uncertainty in high noise or constrained data environments remains insufficient. In situations when there isn’t much data, models can be too sure of themselves or not well-calibrated at all, which makes their outputs less reliable. The current paper presents a theoretical framework elucidating the interaction of uncertainty with data quantity and interpretability in sparse-data contexts. This framework establishes a formal foundation for the development of medical imaging approaches that are both reliable and transparent.
II-D Comparative Theoretical Context
Classical learning theory, from Vapnik’s The Nature of Statistical Learning Theory [20] and McAllester’s PAC-Bayesian theorems [21] to the information-theoretic analyses of Tishby et al. [22] and Xu & Raginsky [23], has provided generalization guarantees under ideal i.i.d. and fully supervised settings. Subsequent works such as Catoni [24] and Dziugaite & Roy [25] refined PAC-Bayesian bounds for deep neural networks, while Achille & Soatto [26] and Russo & Zou [27] examined information compression, invariance, and bias control. Building on these foundations, our framework extends PAC/VC theory to the data-scarce, multimodal, and interpretable medical imaging setting by introducing a synergy term $\Delta_{\mathrm{mm}}$ for multimodal information gain, deriving PAC-Bayesian uncertainty bounds for sparse supervision, and establishing a formal link between explanation stability, model capacity, and sample complexity.
III Problem Formulation
In order to establish formal guarantees on sample complexity, uncertainty, and interpretability, we begin by defining the learning setup, notation, and assumptions used throughout this work.
III-A Notation and Setup
Let each data sample be a tuple $(x,t,y)$, where:
  • $x\in\mathcal{X}$ denotes imaging data (e.g., MRI, CT, histopathology etc.),
  • $t\in\mathcal{T}$ represents complementary structured information such as electronic health records (EHR) or clinical metadata,
  • $y\in\mathcal{Y}$ is the clinical label, which may be categorical (diagnosis) or continuous (severity score).
Let $D_{L}=\{(x_{i},t_{i},y_{i})\}_{i=1}^{n_{L}}\subset\mathcal{X}\times\mathcal{T}\times\mathcal{Y}$ denote the labeled dataset, assumed to be drawn independently and identically distributed (i.i.d.) from an unknown joint distribution $P(x,t,y)$. Let $D_{U}$ represent any available unlabeled or auxiliary data. We consider a hypothesis class $\mathcal{F}$ consisting of predictive functions $f_{\theta}:\mathcal{X}\times\mathcal{T}\rightarrow\mathcal{Y}$ parameterized by $\theta\in\Theta$.
III-B Learning Objective
The model is trained to minimize the expected prediction error over the joint data distribution $(x,t,y)$:
$\mathcal{R}(\theta)=\mathbb{E}_{(x,t,y)}\big[(f_{\theta}(x,t)-y)^{2}\big],$ (1)
where $\mathcal{R}(\theta)$ denotes the expected risk, and $\ell(f_{\theta}(x,t),y)=(f_{\theta}(x,t)-y)^{2}$ represents the squared loss, quantifying the deviation between the model’s prediction and the true label. Depending on the task, this loss can be adapted for classification, regression, or segmentation.
In low-resource regimes, the number of labeled samples $n_{L}$ satisfies $n_{L}\ll N$, where $N$ is the typical sample size required for standard generalization. The objective is to find the smallest number of labeled samples, $n_{L}$, that ensures the model’s expected risk is close to the optimal value:
$\Pr\big[\mathcal{R}(\theta)-\mathcal{R}^{*}\leq\epsilon\big]\geq 1-\delta,$ (2)
where $\mathcal{R}^{*}=\min_{\theta\in\Theta}\mathcal{R}(\theta)$ denotes the lowest attainable risk within the hypothesis space $\Theta$.
III-C Assumptions
To enable theoretical analysis, we adopt the following:
  1. 1.
    Limited Labeled Data: $|D_{L}|=n_{L}\ll N$, reflecting low-resource scenarios.
  2. 2.
    Complementary Modalities: The mutual information between $x$ and $t$ shows that the two modalities encode related but not identical aspects of the data. Since
    I(x;t)<H(x),H(t),I(x;t)<H(x),\;H(t),
    the dependence between them is only partial, which indicates that while they do share some information, each modality also contributes unique, non-overlapping features to the learning process.
  3. 3.
    Label Noise: We model the observed labels as
    $y=y^{*}+\eta,$
    where $y^{*}$ is the true label and $\eta$ represents bounded noise.
  4. 4.
    Hypothesis Class Capacity: The model class $\mathcal{F}$ has finite VC-dimension $\text{VC}(\mathcal{F})$ or bounded Rademacher complexity $\mathfrak{R}_{n}(\mathcal{F})$.
  5. 5.
    i.i.d. Sampling: We consider independent and identically distributed (i.i.d.) sampling, where each sample is drawn from the same underlying distribution and is statistically independent of the others. In particular, the training examples $(x_{i},t_{i},y_{i})$ are drawn i.i.d. from $P$.
Figure 1: Architecture of Low-Resource Learning in Medical Imaging
III-D Problem Scope
Under this framework, we aim to provide rigorous bounds on:
  1. 1.
    Sample Complexity: Minimum labeled data $n_{L}$ required to guarantee $\epsilon$-accurate predictions with confidence $1-\delta$.
  2. 2.
    Multimodal Generalization: How complementary modalities $t$ reduce effective sample complexity or variance.
  3. 3.
    Uncertainty Quantification: Upper bounds on predictive variance or confidence intervals under sparse supervision.
  4. 4.
    Interpretability Guarantees: Stability of explanation methods quantified as a function of sample size and model complexity.
Figure 1 presents the end-to-end pipeline for learning in low-resource conditions. The process begins by feeding both abundant and limited labeled datasets into a learning framework that integrates few-shot samples within a theoretical model $f_{\theta}$. The model is built to not only deliver diagnostic predictions, but also report how confident those predictions are and explain the reasoning behind them. It brings together different types of information, for example medical images and patient records, and learns a joint representation that supports clinical decision making. Uncertainty estimation and explanation are treated as separate analytical paths within the system, each produces an output that is later combined with the core prediction through an aggregation layer. This setup helps the model remain reliable and understandable, even when the amount of labeled training data is small.
III-E Analytical Roadmap
Section IV develops the theoretical bounds in three steps. We begin by examining the sample complexity when multiple complementary modalities are used, and introduce a synergy term, $\Delta_{\text{mm}}$, to describe the information gained by combining them. We define the multimodal synergy term as
$\Delta_{\text{mm}}=I(y;t\mid x)=I((x,t);y)-I(x;y),$
which measures the additional predictive information contributed by modality $t$ beyond $x$. A positive $\Delta_{\text{mm}}$ indicates that the second modality provides complementary, non-redundant information that reduces the required number of labeled samples for a fixed target accuracy. However, $\Delta_{\text{mm}}>0$ does not guarantee improved generalization in every setting; the gain depends on whether the additional modality contributes unique predictive information that aligns with the target variable and is not confounded by noise or redundancy. Second, we obtain uncertainty guarantees via a structured PAC-Bayesian prior that couples modality-specific parameters. Third, we bound explanation variance by leveraging Lipschitz smoothness of the explanation functional and parameter concentration properties from empirical risk minimization and PAC-Bayesian theory.
Section V then translates these mathematical results into deployment guidelines, including label-budget thresholds, confidence-based decision gating, and explanation-stability monitoring.
IV Theoretical Analysis
This section formalizes the mathematical foundations of the proposed framework, linking sample complexity, uncertainty quantification, and interpretability stability within a unified view of low-resource multimodal medical imaging. We derive formal bounds for learning under limited supervision and show how multimodal information and sequential reasoning influence generalization and explanation consistency.
IV-A Sample Complexity and Few-Shot Learning
A fundamental challenge in low-resource learning is determining the minimal number of labeled examples needed to reach clinically acceptable accuracy. In classical PAC-learning theory, the number of labeled samples $n_{L}$ needed to guarantee an expected risk within $\epsilon$ of the optimal value $R^{\star}$ with confidence $1-\delta$ scales as
$n_{L}\geq\frac{C}{\varepsilon^{2}}\left(VC(F)\log\frac{1}{\varepsilon}+\log\frac{1}{\delta}\right),$ (3)
where $VC(\mathcal{F})$ denotes the capacity of the hypothesis class.
Theorem 1 (PAC Sample Complexity:).
Let $\mathcal{F}$ be a hypothesis class with VC-dimension $VC(\mathcal{F})$. For a loss function bounded in $[0,1]$, to achieve $R(\hat{f})-R^{\star}\leq\epsilon$ with probability at least $1-\delta$, it suffices that … it suffices that
$n_{L}\geq\frac{C}{\varepsilon^{2}}\left(VC(F)\log\frac{1}{\varepsilon}+\log\frac{1}{\delta}\right).$
Here, the $\varepsilon^{-2}$ dependence reflects the standard agnostic PAC bound for bounded or sub-Gaussian losses.
Proof sketch.
Follows from uniform convergence and the Sauer–Shelah lemma under i.i.d. sampling. Here, the $\varepsilon^{-2}$ dependence reflects the standard agnostic PAC bound for bounded or sub-Gaussian losses. The constant $C$ absorbs logarithmic and variance terms. For unbounded losses such as squared error, assume sub-Gaussian noise or apply a clipped surrogate to ensure bounded variance. Throughout, we normalize all losses to lie in $[0,1]$ (by scaling or clipping) to satisfy PAC and PAC–Bayesian bounded-loss assumptions.
When complementary modalities such as imaging $x$ and structured clinical data $t$ are available, a useful model family is $\mathcal{F}_{x,t}=\{\,f(x,t)=g(x)+h(t):g\in\mathcal{F}_{x},\;h\in\mathcal{F}_{t}\,\}$. The combined capacity then satisfies the sub-additive property:
Proposition 1 (Sub-additive Pseudo-dimension for Multimodal Models).
For binary classification with a thresholded linear combiner $f(x,t)=\mathrm{sign}(g(x)+h(t))$, or for real-valued predictors under pseudo-dimension analysis, one has
$\mathrm{Pdim}(\mathcal{F}_{x,t})\;\leq\;\mathrm{Pdim}(\mathcal{F}_{x})\;+\;\mathrm{Pdim}(\mathcal{F}_{t}).$
Proof sketch. For classification, this follows from the sub-additivity of the growth function under summation of hypothesis classes; for regression, the analogous inequality holds for the pseudo-dimension by extending the argument to real-valued outputs.
This implies that multimodal learning can reduce the effective data requirement by leveraging shared but non-redundant information between modalities.
Under an $N$-way $K$-shot setting, the expected generalization error scales as $O(1/\sqrt{m})$ with $m=NK$ i.i.d. labeled samples, assuming tasks and examples are drawn independently, consistent with meta-learning analyses showing that modest increases in per-class supervision can yield substantial gains.
IV-B Uncertainty Quantification via PAC–Bayes Bounds
Reliable clinical systems must not only be accurate but also quantify predictive confidence. The predictive variance can be written as
$\mathrm{Var}[Y\,|\,x,t]\;=\;\int(y-\mathbb{E}[Y\,|\,x,t])^{2}\,p(y\,|\,x,t)\,dy,$ (4)
which measures the dispersion of outcomes given inputs $(x,t)$. Within the PAC–Bayesian framework, the expected risk of a stochastic model with posterior $Q$ and prior $P$ satisfies:
Theorem 2 (PAC–Bayesian Risk Bound).
With probability at least $1-\delta$ over $n_{L}$ i.i.d. samples, for any prior $P$ and posterior $Q$,
$\mathbb{E}_{\theta\sim Q}\!\big[L(\theta)\big]\;\leq\;\hat{L}_{Q}\;+\;\sqrt{\frac{KL(Q\|P)+\ln(1/\delta)}{2\,n_{L}}}\,$,
where $\hat{L}_{Q}=\mathbb{E}_{\theta\sim Q}\!\big[\hat{L}(\theta)\big]$ is the empirical loss. Proof sketch. Follows from McAllester’s PAC–Bayesian theorem using change of measure and exponential concentration.
When multiple correlated modalities constrain the parameter space, the divergence term $KL(Q\|P)$ can decrease, tightening the bound and yielding lower predictive uncertainty.
IV-C Interpretability and Explanation Stability
Interpretability requires that explanations remain consistent under small perturbations in data or model parameters. Let $E(f_{\theta},x,t)$ denote the explanation functional (e.g., a feature attribution or saliency value at a fixed location).
Assumption 1 (Lipschitz Regularity).
The explanation map is $L$-Lipschitz in model parameters:
$\big|E(f_{\theta_{1}},x,t)\;-\;E(f_{\theta_{2}},x,t)\big|\;\leq\;L\,\|\theta_{1}-\theta_{2}\|\quad\text{for all }\theta_{1},\theta_{2}.$
Theorem 3 (Explanation Variance Bound).
Under Assumption 1, i.i.d. samples, and a hypothesis class $\mathcal{F}$ with finite $VC(\mathcal{F})$, the variance of explanations satisfies
$\mathrm{Var}\!\big[E(f_{\theta},x,t)\big]\;\leq\;C\,\frac{VC(\mathcal{F})}{n_{L}},$
for a constant $C$ depending on $L$ and the loss range. Proof sketch. Parameter concentration around an empirical minimizer occurs at rate $\mathcal{O}\!\big(\sqrt{VC(\mathcal{F})/n_{L}}\big)$ by uniform convergence or PAC–Bayes. Lipschitz continuity transfers this concentration to explanation outputs, yielding the inverse-$n_{L}$ scaling.
As $n_{L}$ increases or models are better regularized, explanations become more stable, providing a quantitative basis for interpretability guarantees.
IV-D Sequential Reasoning and Posterior Contraction
The proposed Chain-of-Thought (CoT) reasoning can be interpreted as sequential Bayesian updates:
$p(y\,|\,s_{i},x,t)\;=\;\frac{p(s_{i}\,|\,y,x,t)}{p(s_{i}\,|\,x,t)}\;p(y\,|\,s_{i-1},x,t),$ (5)
where each step incorporates additional evidence $s_{i}$ that refines the belief over $y$.
Claim 1 (Stepwise Posterior Contraction).
Let $Q_{i}$ denote the posterior distribution over $\theta$ after step $i$. If each $s_{i}$ provides conditionally independent evidence about $y$ given prior steps, then
$\mathbb{E}\!\left[\mathrm{KL}(Q_{i}\,\|\,P)\right]=\mathbb{E}\!\left[\mathrm{KL}(Q_{i-1}\,\|\,P)\right]+I(\theta;s_{i}\mid x,t,s_{<i}).$
Interpretation. Each reasoning step contributes a non-negative information gain $I(\theta;s_{i}\mid x,t,s_{<i})$ that refines the posterior. Contraction occurs not in $\mathrm{KL}(Q_{i}\|P)$ itself but in the posterior entropy $H(Q_{i})$ or in its divergence to the true parameter distribution.
Consequently, both uncertainty and explanation variance contract across reasoning steps, linking the CoT process to the theoretical quantities introduced above.
IV-E Trade-Offs and Insights
The derived results lead to several practical observations:
  • Accuracy vs. Data. Larger models can achieve higher accuracy but require more labeled data to maintain generalization.
  • Uncertainty vs. Complexity. Multimodal data can mitigate overfitting and improve confidence estimates by regularizing the posterior through shared evidence.
  • Interpretability vs. Robustness. Explanation stability improves with sample size and regularization, supporting model auditing in low-resource settings.
Together, Theorems 13 and Claim 1 establish a unified foundation for low-resource multimodal learning with uncertainty-aware explainability.
V Implications for Real-World Deployment
The analysis above offers several practical lessons for using AI in low-resource medical imaging settings.
V-A Data and Model Requirements
Sample complexity bounds describe how much labeled data are needed to achieve reliable accuracy. If the model’s hypothesis class has VC-dimension $\text{VC}(\mathcal{F})$, then for a target error $\epsilon$ and confidence level $1-\delta$:
$n_{L}\geq\frac{C}{\varepsilon^{2}}\left(VC(F)\log\frac{1}{\varepsilon}+\log\frac{1}{\delta}\right)$ (6)
which gives a way to estimate whether the dataset is large enough for the task.
When different data modalities are used together, the effective capacity of the combined model is smaller:
$\text{VC}(\mathcal{F}_{x,t})\leq\text{VC}(\mathcal{F}_{x})+\text{VC}(\mathcal{F}_{t}),$ (7)
indicating that multimodal learning can reduce data needs and improve robustness when labeled samples are limited.
V-B Uncertainty- and Explanation-Aware Deployment
PAC–Bayesian analysis provides a principled framework for making confidence-aware decisions. For a posterior distribution $Q$ over model parameters $\theta$, the expected risk satisfies the bound:
$\mathbb{E}_{\theta\sim Q}\!\left[\mathcal{L}(\theta)\right]\leq\hat{\mathcal{L}}_{Q}+\sqrt{\frac{\mathrm{KL}(Q\,\|\,P)+\ln\!\left(\tfrac{1}{\delta}\right)}{2n_{L}}},$ (8)
where $\hat{\mathcal{L}}_{Q}=\mathbb{E}_{\theta\sim Q}[\hat{\mathcal{L}}(\theta)]$ is the expected empirical loss under the posterior $Q$. The term $P$ denotes the prior distribution, and $\mathrm{KL}(Q\,\|\,P)$ measures how far the posterior departs from the prior. This yields a probabilistic link between empirical performance and its expected generalization.
A similar idea applies to the behaviour of explanations. In particular,
$\mathrm{Var}[E(f_{\theta},x,t)]\leq\mathcal{O}\!\left(\frac{\mathrm{VC}(\mathcal{F})}{n_{L}}\right),$ (9)
which shows that explanation variability decreases when more labeled data are available or when the model class is less complex. This is important in clinical settings, where explanations need to be steady and trustworthy.
V-C Deployment Guidelines
  • Use multimodal data whenever possible to reduce labeling requirements and increase robustness.
  • Set confidence-based thresholds to initiate expert review in uncertain instances.
  • Match model capacity with data size to maintain accuracy, uncertainty, and interpretability, and monitor explanation stability on frequently, especially with the training data is low.
VI Open Theoretical Problems
Despite the theoretical bounds presented in this work, several challenges remain. Multimodal integration appears to lower data needs in practice, but its information-theoretic basis is still unclear. Important questions include how mutual information between modalities affects generalization and which combinations of modalities are enough to achieve reliable accuracy when data are limited. Most current analyses also assume i.i.d. sampling, which rarely holds in clinical data that vary across sites, equipment, and patient groups.
At the same time, interpretability and robustness need stronger theoretical support. Current widely used explanation methods provide few guarantees when training data are scarce, and the trade-offs between accuracy, uncertainty, and interpretability are not yet well understood. So, future work should focus on building information-theoretic models that connect these aspects and provide formal robustness guarantees, helping make clinical AI both efficient and dependable.
VII Conclusion
This paper presents a theoretical framework for low-resource medical imaging that brings together sample complexity, uncertainty quantification, and interpretability stability within a single formal setting. Using concepts from PAC learning, VC-dimension theory, and PAC-Bayesian analysis, we establish bounds for:
  • Sample Complexity: The least number of labeled samples needed to ensure clinically valid model accuracy.
  • Multimodal Learnability: How combining complementary data sources lowers the effective data requirement.
  • Uncertainty Quantification: Limits on predictive variance when working with small or noisy datasets.
  • Interpretability Guarantees: How explanation stability depends on both, data availability and model complexity.
Through a principled theoretical foundation, our approach advances medical imaging toward models capable of safe, interpretable, and effective clinical deployment.